# Compressing multivariate functions with tree tensor networks

### v1

@@ I Introduction
Tensor trains also offer a somewhat unconventional numerical methodology for problems in continuous space [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Using an encoding of the relevant continuous variables into binary strings of length $L$, mathematical functions on a grid with spacing $\mathcal{O}(\exp(-L))$ can be represented with a train of $\mathcal{O}(L)$ order-three tensors. Such an ansatz is commonly referred to as a quantics tensor train (QTT) and has opened up a new field of tensor train-based numerical methods. Whilst tensor trains are known to be highly effective for smooth, one-dimensional functions [31], higher-dimensional functions can pose significant difficulties, typically requiring much larger ranks for the tensors in the train and thus larger computational resources. Other than mathematical and computational simplicity, however, there is no reason to limit these tensor-based numerical methods to trains. Tensor networks of more complex topology—which have proven fundamental in the field of two-dimensional quantum simulation [32, 33, 34, 35]—offer a whole new degree of freedom, allowing more structured, complex correlations to be encoded between the underlying variables (which, in this context, are the binary digits). Several works have considered ansatzes beyond the tensor train for representing multivariate functions. Specifically: multiple tensor trains coupled via their leading tensors [36, 37] or “functional” tensor trains where the individual tensors in the train constitute matrix-valued functions [38, 39]. Very few methods are available in this domain, however, for working with tensor networks of more generic topology and little is understood about how their structure determines their effectiveness at representing a given continuous function.
@@ II Preliminaries
We define a tensor network as a connected network of tensors: each vertex $v$ of the network hosts a tensor and the edges of the network dictate which tensors share common ‘virtual’, ‘internal’, or ‘bond’ indices. The maximum dimension of any of the virtual (or bond) indices in the network is referred as the bond dimension or rank $\chi$ of the tensor network. Each tensor of a tensor network can also have external indices not common to any other tensors in the network. A tensor network with a total of $N$ external indices corresponds to a decomposition of an order $N$ tensor.
@@
In this work the external indices of the network represent discrete variables which decompose a series of $n$ continuous variables $\mathbf{x}=(x_{1},x_{2},...,x_{n})\in[0,1)^{n}$ in a binary manner. Specifically, for a given continuous variable $x_{i}$ this binary decomposition reads $x_{i}=\sum_{j=1}^{L}\frac{x_{i,j}}{2^{j}}$ where the $x_{i,j}\in\{0,1\}$ are the binary variables or bits which are each represented by an external index of dimension $2$ in the tensor network  [^1] In this work, for simplicity, we assume binary strings which are all of length $L$ and thus the total number of external indices is $nL$. The generalization of our work to $n-$ary decompositions with a variable number of binary digits for each variable is straightforward.. The $2^{nL}$ possible binary strings, or configurations of the bits, realises a uniform discrete grid for $\mathbf{x}\in[0,1-\delta]^{n}$ where $\delta=2^{-L}$ is the grid spacing in each dimension which is exponentially ‘fine’ in the number of bits. Unless specified otherwise we will focus on the case where each tensor in the network contains one external index and therefore corresponds to one binary digit $x_{i,j}$ in the decomposition of a continuous variable $x_{i}$. We will focus exclusively on tensor networks which are trees, i.e. networks where the virtual indices do not form loops. This means that they can be optimised and contracted efficiently and we will frequently refer to them as tree tensor networks (TTNs).
@@
Figure 1: a-c) Representations of the two dimensional function $f(\mathbf{x})=f(x_{1},x_{2})$ with the continuous variables $x_{1},x_{2}\in[0,1)$ encoded as binary strings $x_{1}=0.x_{1,1}x_{1,2}x_{1,3}x_{1,4}$ and $x_{2}=0.x_{2,1}x_{2,2}x_{2,3}x_{2,4}$ of length four. a) The actual values of the function over the domain specified by the binary digits can be encoded as a single order $8$ tensor. b) Quantics tensor train or Matrix Product State representation of the order $8$ tensor with a binary digit ordering commonly referred to as ‘sequential’. c) An example of a tree tensor network (TTN) representation of the order $8$ tensor. d) Top: We use the notation $\mathcal{T}_{(i,j)}$ to refer to the local tensor in a tree corresponding to the binary digit $x_{i,j}$ – the $j$th most significant binary digit in the decomposition of the $i$th continuous variable $x_{i}$. This tensor has $z_{(i,j)}$ ‘virtual’ indices (black lines) connecting it to its neighbors in the tree and a single external index (dotted line) corresponding to the binary variable $x_{(i,j)}\in\{0,1\}$. Bottom: We use $\mathcal{T}_{(i,j)}(x_{i,j})$ to refer to the order $z_{(i,j)}$ tensor which is a slice of $\mathcal{T}_{(i,j)}$ for the given value of $x_{i,j}$.
@@
The TTNs in this paper have a structure specified by a labelled tree $\mathcal{T}$ where each of the vertices corresponds to a single binary digit $x_{i,j}$ in the decomposition of $x_{i}$. We will use the notation $\mathcal{T}_{(i,j)}$ to denote the tensor on a given vertex and $\mathcal{T}_{(i,j)}(x_{i,j})$ to refer to a given ‘slice’ of that tensor $\mathcal{T}_{(i,j)}$ for a specific value of $x_{i,j}\in\{0,1\}$, effectively viewing it as a ‘function’ of the binary variable $x_{i,j}$. Each $\mathcal{T}_{(i,j)}(x_{i,j})$ is just another tensor and has order $z_{i,j}$: the co-ordination number of the given vertex $x_{i,j}$ in the tensor network. This idea is illustrated in Fig. 1d. We refer to the $z_{i,j}$ indices connecting a local tensor to its neighbors as ‘internal’ or ‘virtual’ indices. We frequently use the notation $\bm{\alpha}_{i,j}=(\alpha_{1},\alpha_{2},\ldots\alpha_{z_{i,j}})$ to denote the virtual indices on the tensor $\mathcal{T}_{(i,j)}$ and index them starting from 00, i.e. $\alpha_{k}=0,1,2,\ldots,{\rm dim}(\alpha_{k})$.
@@ IV Numerical Results for Function Construction
Figure 3: Comparison of different tree tensor networks, with $L=16$ vertices, for compressing two different one-dimensional functions $f(x)$ with a binary decomposition $x=0.x_{1}x_{2}...x_{16}$. Error $\epsilon$ is calculated via Eq.(5), sampling the function over $10^{3}$ random grid points. The different labelled trees used are shown at the top and the bits are numbered from most significant to least significant: with the most significant digit circled in red. Two functions are considered. Top plots) the Laguerre polynomial $f(x)=L_{n}(x)=\sum_{k=0}^{n}\binom{n}{k}\frac{(-1)^{k}}{k!}x^{k}$ with $n=40$ and Bottom plots) the Weierstrass function $f(x)=\sum_{k=1}^{n}\frac{\sin(\pi k^{a}x)}{\pi k^{a}}$ with $a=3$ and $n=25$. Left panels: Sketch of the functions considered over $x\in[0,1]$. Inset of d) shows a zoomed-in region of the function with data points corresponding to the values obtained from the Bident tree tensor network with $\chi=36$. Middle panels: Error versus bond dimension of the tensor networks. Right panels: Error versus memory requirement for storing the tensor networks —assuming $8$ bytes for a floating point number.
@@
Figure 4: Comparison of different tree tensor networks for compressing the three dimensional function $f(\mathbf{r})=\sum_{j=1}^{n}\cos(j\mathbf{k_{j}}\cdot\mathbf{r})$ with $n=30$, $\mathbf{r}=(x,y,z)$ and $\mathbf{k_{j}}=(k^{x}_{j},k^{y}_{j},k^{z}_{j})$ with $k^{\alpha}_{j}\in\mathcal{N}(0,1)$. Error is calculated as $\epsilon$ via Eq.(5), sampling the function over $10^{3}$ random grid points. Different trees used are shown at the top. Bits for $x$, $y$ and $z$ are coloured in green, grey and light blue respectively and numbered sequentially from most significant to least significant. a) Error versus bond dimension. Inset shows a heatmap of the function at $z=\frac{1}{2}$. b) Error versus memory requirement for storing the tensor networks — assuming $8$ bytes for a single floating point number. c) Mutual information matrix encoding the correlations between the binary digits. Calculated by sampling the function $10^{4}$ times to build up an approximate reduced density matrix for the given paur of bits. Each dashed block encodes the matrix of correlations between the bits of two given dimensions. d) Absolute error over one-dimensional slices of the function. Bar chart shows memory requirements on the right for the specified TTN with the maximum bond dimension of the network also shown. Inset (black line) shows a slice of the function for $y=1/2$ and $z=1/2$, with corresponding results for the coupled binary tree tensor network at $\chi=18$.
@@
3D Function Construction - In Fig. (4 we move on to consider a three-dimensional function which is the sum of $n=30$ random plane waves of increasing frequency and compare three topologies: two tensor trains with different, commonly used, digit orderings (interleaved and sequential) and a tree consisting of three separate binary trees, one for each dimension, coupled together at their roots (where the most significant bits are placed). Here we find that this coupled binary tree tensor network (BTTN) is a better ansatz for the function at hand and compresses much more effectively under truncation of the internal bonds. For a fixed memory cost it can achieve orders of magnitude lower error than the trains. For instance, it is able to represent the function with an error $\epsilon\sim 10^{-6}$ with a bond dimension of $\chi=15$ whilst the tensor trains each require a bond dimension of $\chi=30$ to reach such an error. As $n=30$ the function can be represented exactly on any network with a bond dimension $\chi=30$ and thus this shows that the tensor trains are an ineffective representation. In Fig. 4d) we show the absolute error over multiple one-dimensional slices of the function of the tree TTNs for fixed bond dimensions. Despite having a slightly lower memory cost than the tensor trains, the error in the BTTN is consistently several orders of magnitude lower than them.
@@
Figure 5: Comparison of the effectiveness of different tree tensor networks with $L=16$ bits per dimension for learning the multinormal probability density function $f(\mathbf{r})\propto\exp(-((\mathbf{r}-\bm{\mu})^{T}M^{-1}(\mathbf{r}-\bm{\mu}))$ via the tensor cross interpolation algorithm. Here $M$ is an $n\times n$ covariance matrix and $\bm{\mu}=(\mu_{1},\mu_{2},...,\mu_{n})$ is the mean vector. We consider $n=3$ with $\mathbf{r}=(x,y,z)\in[0,10)^{3}$ and $\mathbf{\mu}=(5,5,5)$. Results are obtained from drawing $N=10$ instances of $M$ ($M_{1},M_{2},\ldots M_{10}$) from the LKJ distribution [51] with shape parameter $\eta=50$. Different trees used are shown at the top. Bits for $x$, $y$ and $z$ are coloured in green, grey and light blue respectively. The most significant digit in each dimension is circled in red. a-b) Error $\epsilon$, calculated via Eq. (5) after $n=10$ sweeps of the TCI algorithm, versus bond dimension and memory cost for the tensor networks. The solid lines shows the mode of the error over the $10$ realizations of $M$ whilst the shaded area shows the range of the error, i.e. for any $M$ sampled the error $\epsilon$ for a given bond dimension. Inset) shows a heatmap of the function over $(x,z)\in[3,7]^{2}$ with $y=\frac{1}{2}$. c) Average value for the infinity norm $\epsilon_{\infty}$ (see Eq. (5)) over a given sweep of the TCI algorithm for $M=M_{1}$ and the three tensor networks at the specified bond dimensions and memory cost.
TCI Function Construction - In Fig. 5 we use the TCI algorithm to build representations of the trivariate gaussian probability density function $f(\mathbf{r})\propto\exp(-(\mathbf{r}-\bm{\mu})^{T}M^{-1}(\mathbf{r}-\bm{\mu}))$ where $\mathbf{r}=(x,y,z)$, $\bm{\mu}=(\mu_{x},\mu_{y},\mu_{z})$ is the mean vector, and $M$ is a covariance matrix which we sample from the Lewandowski-Kurowicka-Joe (LKJ) distribution with shape parameter $\eta=50$ [51], which controls the weight of the off-diagonal correlations. We compare results from the TCI algorithm when varying the maximum allowed bond dimension for different tree tensor networks: two tensor trains with commonly used digit orderings (sequential and interleaved) for multi-dimensional functions and a comb tree tensor network consisting of sequentially ordered tensor trains (CTTN - Seq) coupled by their most significant digit.

### v2

@@ I Introduction
Tensor trains also offer a somewhat unconventional numerical methodology for problems in continuous space [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Using an encoding of the relevant continuous variables into binary strings of length $L$, mathematical functions on a grid with spacing $\mathcal{O}(\exp(-L))$ can be represented with a train of $\mathcal{O}(L)$ order-three tensors. Such an ansatz is commonly referred to as a quantics tensor train (QTT) and has opened up a new field of tensor train-based numerical methods. Whilst tensor trains are known to be highly effective for smooth, one-dimensional functions [32], higher-dimensional functions can pose significant difficulties, typically requiring much larger ranks for the tensors in the train and thus larger computational resources. Other than mathematical and computational simplicity, however, there is no reason to limit these tensor-based numerical methods to trains. Tensor networks of more complex topology—which have proven fundamental in the field of two- and three-dimensional quantum simulation [33, 34, 35, 36, 37, 38]—offer a whole new degree of freedom, allowing more structured, complex correlations to be encoded between the underlying variables (which, in this context, are the binary digits). Several works have considered ansatzes beyond the tensor train for representing multivariate functions. Specifically: multiple tensor trains coupled via their leading tensors [39, 40], “functional” tensor trains where the individual tensors in the train constitute matrix-valued functions [41, 42] or heirarchical tucker decompositions where the binary variables are placed solely at the leaves of the tree and correlated with each other via their parents [43, 44]. Very few methods are available in this domain, however, for working with tensor networks of more generic topology and little is understood about how their structure determines their effectiveness at representing a given continuous function.
@@ II Preliminaries
A tensor $\mathcal{T}_{v}$ is a function which maps discrete variables (which we refer to as its indices) to numbers, i.e. $\mathcal{T}_{v}(a_{1},a_{2},\ldots)\in\mathbb{C}$ with $a_{i}\in[1,2,\ldots,d_{i}]$. A tensor network is an interconnected network of tensors: each vertex of the network $v\in V$, with $V$ the set of vertices of the network, hosts a tensor $\mathcal{T}_{v}$ and the edges of the network dictate which tensors share common ‘virtual’, ‘internal’, or ‘bond’ indices that are implicitly summed over. The maximum dimension of any of the virtual (or bond) indices in the network is referred as the bond dimension or rank $\chi$ of the tensor network. Each tensor of a tensor network can also have external indices not common to any other tensors in the network. Defining $\mathbf{I}$ as the set of all indices in the network, $\bm{\alpha}$ as the set of all indices which appear twice (i.e. are shared by two tensors) and $\mathbf{x}=\mathbf{I}\setminus\bm{\alpha}$ as the set of all external indices we have
$\mathcal{T}(\mathbf{x})=\sum_{\bm{\alpha}}\prod_{v\in V}\mathcal{T}_{v}(\mathbf{I_{v}}),\qquad\mathbf{I_{v}},\bm{\alpha}\subseteq\mathbf{x},$ (1)
with $\mathbf{I_{v}}$ the indices possesed by the tensor $\mathcal{T}_{v}.$ The tensor network $\mathcal{T}$ thus corresponds to a factorization of a single tensor with indices $\mathbf{x}$. Some example tensor networks are illustrated in Fig. 1b) and c) along with the tensor they are representing in Fig. 1a).
@@
In this work the external indices of the network represent discrete variables which decompose a series of $n$ continuous variables $\mathbf{x}=(x_{1},x_{2},...,x_{n})\in[0,1)^{n}$ in a binary manner. For a given continuous variable $x_{i}$ the binary decomposition reads $x_{i}=\sum_{j=1}^{L}\frac{x_{i,j}}{2^{j}}$ where the $x_{i,j}\in\{0,1\}$ are the binary variables or bits which are each represented by an external index of dimension $2$ in the tensor network  [^1] In this work, for simplicity, we assume binary strings which are all of length $L$ and thus the total number of external indices is $nL$. The generalization of our work to $n$-ary decompositions with a variable number of binary digits for each variable is straightforward.. The $2^{nL}$ possible binary strings, or configurations of the bits, realises a uniform discrete grid for $\mathbf{x}\in[0,1-\delta]^{n}$ where $\delta=2^{-L}$ is the grid spacing in each dimension which is exponentially ‘fine’ in the number of bits. Unless specified otherwise we will focus on the case where each tensor in the network contains one external index and therefore corresponds to one binary digit $x_{i,j}$ in the decomposition of a continuous variable $x_{i}$. We will focus exclusively on tensor networks which are trees, i.e. networks where the virtual indices do not form loops. We should emphasize that this is related, but not the same as the previously introduced hierarchical tucker format [43, 44, 48] where the binary variables live only on the leaves of the tensor network and tensors in the bulk are simply factors that mediate correlations between them and do not contain the external variables. The fact the tensor networks we consider do not contain loops this means that they can be optimised and contracted (for a given configuration of their binary variables) efficiently and we will refer to them as tree tensor networks (TTNs).
@@
Figure 1: a-c) Representations of the two dimensional function $f(\mathbf{x})=f(x_{1},x_{2})$ with the continuous variables $x_{1},x_{2}\in[0,1)$ encoded as binary strings $x_{1}=0.x_{1,1}x_{1,2}x_{1,3}x_{1,4}$ and $x_{2}=0.x_{2,1}x_{2,2}x_{2,3}x_{2,4}$ of length four. a) The actual values of the function over the domain specified by the binary digits can be encoded as a single order 88 tensor. b) Quantics tensor train or Matrix Product State representation of the order 88 tensor with a binary digit ordering commonly referred to as ‘sequential’. c) An example of a tree tensor network (TTN) representation of the order 88 tensor. d) Top: We use the notation $\mathcal{T}_{(i,j)}$ to refer to the local tensor in a tree corresponding to the binary digit $x_{i,j}$ – the $j$th most significant binary digit in the decomposition of the $i$th continuous variable $x_{i}$. This tensor has $z_{(i,j)}$ ‘virtual’ indices (black lines) connecting it to its neighbors in the tree and a single external index (dotted line) corresponding to the binary variable $x_{(i,j)}\in\{0,1\}$. Bottom: We use $\mathcal{T}_{(i,j)}(x_{i,j})$ to refer to the order $z_{(i,j)}$ tensor which is a slice of $\mathcal{T}_{(i,j)}$ for the given value of $x_{i,j}$.
@@
The TTNs in this paper have a structure specified by a labelled tree $\mathcal{T}$ where each of the vertices corresponds to a single binary digit $x_{i,j}$ in the decomposition of $x_{i}$. We will use the notation $\mathcal{T}_{(i,j)}$ to denote the tensor on a given vertex and $\mathcal{T}_{(i,j)}(x_{i,j})$ to refer to a given ‘slice’ of that tensor $\mathcal{T}_{(i,j)}$ for a specific value of $x_{i,j}\in\{0,1\}$, effectively viewing it as a ‘function’ of the binary variable $x_{i,j}$. Each $\mathcal{T}_{(i,j)}(x_{i,j})$ is just another tensor and has order $z_{i,j}$: the co-ordination number of the given vertex $x_{i,j}$ in the tensor network. This idea is illustrated in Fig. 1d. We refer to the $z_{i,j}$ indices connecting a local tensor to its neighbors as ‘internal’ or ‘virtual’ indices. We frequently use the notation $\bm{\alpha}_{i,j}=(\alpha_{1},\alpha_{2},\ldots\alpha_{z_{i,j}})$ to denote the virtual indices on the tensor $\mathcal{T}_{(i,j)}$ and index them starting from 0, i.e. $\alpha_{k}=0,1,2,\ldots,{\rm dim}(\alpha_{k})$.
@@ IV Numerical Results for Function Construction
Figure 3: Comparison of different tree tensor networks, with $L=16$ vertices, for compressing two different one-dimensional functions $f(x)$ with a binary decomposition $x=0.x_{1}x_{2}...x_{16}$. Error $\epsilon$ is calculated via Eq.(6), sampling the function over $10^{3}$ random grid points. The different labelled trees used are shown at the top and the bits are numbered from most significant to least significant: with the most significant digit circled in red. Two functions are considered. Top plots) the Laguerre polynomial $f(x)=L_{n}(x)=\sum_{k=0}^{n}\binom{n}{k}\frac{(-1)^{k}}{k!}x^{k}$ with $n=40$ and Bottom plots) the Weierstrass function $f(x)=\sum_{k=1}^{n}\frac{\sin(\pi k^{a}x)}{\pi k^{a}}$ with $a=3$ and $n=25$. Left panels: Sketch of the functions considered over $x\in[0,1]$. Inset of d) shows a zoomed-in region of the function with data points corresponding to the values obtained from the Bident tree tensor network with $\chi=36$. Middle panels: Error versus bond dimension of the tensor networks. Right panels: Error versus memory requirement for storing the tensor networks —assuming 88 bytes for a floating point number.
@@
Figure 4: Comparison of different tree tensor networks for compressing the three dimensional function $f(\mathbf{r})=\sum_{j=1}^{n}\cos(j\mathbf{k_{j}}\cdot\mathbf{r})$ with $n=30$, $\mathbf{r}=(x,y,z)$ and $\mathbf{k_{j}}=(k^{x}_{j},k^{y}_{j},k^{z}_{j})$ with $k^{\alpha}_{j}\in\mathcal{N}(0,1)$. Error is calculated as $\epsilon$ via Eq.(6), sampling the function over $10^{3}$ random grid points. Different trees used are shown at the top. Bits for $x$, $y$ and $z$ are coloured in green, grey and light blue respectively and numbered sequentially from most significant to least significant. a) Error versus bond dimension. Inset shows a heatmap of the function at $z=\frac{1}{2}$. b) Error versus memory requirement for storing the tensor networks — assuming 88 bytes for a single floating point number. c) Mutual information matrix encoding the correlations between the binary digits. Calculated by sampling the function $10^{4}$ times to build up an approximate reduced density matrix for the given paur of bits. Each dashed block encodes the matrix of correlations between the bits of two given dimensions. d) Absolute error over one-dimensional slices of the function. Bar chart shows memory requirements on the right for the specified TTN with the maximum bond dimension of the network also shown. Inset (black line) shows a slice of the function for $y=1/2$ and $z=1/2$, with corresponding results for the coupled binary tree tensor network at $\chi=18$.
@@
3D Function Construction - In Fig. 4 we move on to consider a three-dimensional function which is the sum of $n=30$ random plane waves of increasing frequency and compare three topologies: two tensor trains with different, commonly used, digit orderings (interleaved and sequential) and a tree consisting of three separate binary trees, one for each dimension, coupled together at their roots (where the most significant bits are placed). Here we find that this coupled binary tree tensor network (BTTN) is a better ansatz for the function at hand and compresses much more effectively under truncation of the internal bonds. For a fixed memory cost it can achieve orders of magnitude lower error than the trains. For instance, it is able to represent the function with an error $\epsilon\sim 10^{-6}$ with a bond dimension of $\chi=15$ whilst the tensor trains each require a bond dimension of $\chi=30$ to reach such an error. As $n=30$ the function can be represented exactly on any network with a bond dimension $\chi=30$ and thus this shows that the tensor trains are an ineffective representation. In Fig. 4d) we show the absolute error over multiple one-dimensional slices of the function of the tree TTNs for fixed bond dimensions. Despite having a slightly lower memory cost than the tensor trains, the error in the BTTN is consistently several orders of magnitude lower than them.
@@
Figure 5: Comparison of the effectiveness of different tree tensor networks with $L=16$ bits per dimension for learning the multinormal probability density function $f(\mathbf{r})\propto\exp(-((\mathbf{r}-\mathbf{\mu})^{T}M^{-1}(\mathbf{r}-\mathbf{\mu}))$ via the tensor cross interpolation algorithm. Here $M$ is an $n\times n$ covariance matrix and $\mathbf{\mu}=(\mu_{1},\mu_{2},...,\mu_{n})$ is the mean vector. We consider $n=3$ with $\mathbf{r}=(x,y,z)\in[0,10)^{3}$ and $\mathbf{\mu}=(5,5,5)$. Results are obtained from drawing $N=10$ instances of $M$ ($M_{1},M_{2},\ldots M_{10}$) from the LKJ distribution [57] with shape parameter $\eta=50$. Different trees used are shown at the top. Bits for $x$, $y$ and $z$ are coloured in green, grey and light blue respectively. The most significant digit in each dimension is circled in red. a-b) Error $\epsilon$, calculated via Eq. (6) after $n=10$ sweeps of the TCI algorithm, versus bond dimension and memory cost for the tensor networks. The solid lines shows the mode of the error over the 1010 realizations of $M$ whilst the shaded area shows the range of the error, i.e. for any $M$ sampled the error $\epsilon$ for a given bond dimension. Inset) shows a heatmap of the function over $(x,z)\in[3,7]^{2}$ with $y=\frac{1}{2}$. c) Average value for the infinity norm $\epsilon_{\infty}$ (see Eq. (6)) over a given sweep of the TCI algorithm for $M=M_{1}$ and the three tensor networks at the specified bond dimensions and memory cost.
TCI Function Construction - In Fig. 5 we use the TCI algorithm to build representations of the trivariate gaussian probability density function $f(\mathbf{r})\propto\exp(-(\mathbf{r}-\mathbf{\mu})^{T}M^{-1}(\mathbf{r}-\mathbf{\mu}))$ where $\mathbf{r}=(x,y,z)$, $\mathbf{\mu}=(\mu_{x},\mu_{y},\mu_{z})$ is the mean vector, and $M$ is a covariance matrix which we sample from the Lewandowski-Kurowicka-Joe (LKJ) distribution with shape parameter $\eta=50$ [57], which controls the weight of the off-diagonal correlations. We compare results from the TCI algorithm when varying the maximum allowed bond dimension for different tree tensor networks: two tensor trains with commonly used digit orderings (sequential and interleaved) for multi-dimensional functions and a comb tree tensor network consisting of sequentially ordered tensor trains (CTTN - Seq) coupled by their most significant digit.